OpenStudy (anonymous):

A student is trying to solve the system of two equations given below: Equation P: y + z = 6 Equation Q: 5y + 9z = 1 Which of the following is a possible step used in eliminating the y-term? (y + z = 6) ⋅ 9 (y + z = 6) ⋅ −5 (5y + 9z = 1) ⋅ 9 (5y + 9z = 1) ⋅ 5

OpenStudy (laughoutloud):

how are the questions setup? When they are together its hard to tell what choices you have to choose from.

OpenStudy (anonymous):

the bottom is the answers

OpenStudy (laughoutloud):

ik but where does each of the answers end?

OpenStudy (anonymous):

@carolinar7

OpenStudy (carolinar7):

hey

OpenStudy (anonymous):

help

OpenStudy (carolinar7):

ok

OpenStudy (carolinar7):

what do you think it is

OpenStudy (anonymous):

c

OpenStudy (carolinar7):

re write the multiple choice question because I cant understand it

OpenStudy (anonymous):

A student is trying to solve the system of two equations given below: Equation P: y + z = 6 Equation Q: 5y + 9z = 1 Which of the following is a possible step used in eliminating the y-term? (y + z = 6) ⋅ 9 (y + z = 6) ⋅ −5 (5y + 9z = 1) ⋅ 9 (5y + 9z = 1) ⋅ 5

OpenStudy (carolinar7):

Is it A. (y+z=6) B.9(y+z)=6 C.-5(5y+9z=1) D.9(5y+9z=1)5

OpenStudy (anonymous):

yes

OpenStudy (carolinar7):

I dont understand it @ganeshie8

OpenStudy (carolinar7):

@pooja195

OpenStudy (carolinar7):

@vera_ewing

OpenStudy (carolinar7):

@ranga

OpenStudy (carolinar7):

@ybarrap